OpenStudy (thatonegirl_):

find the inverse: f(x)=1/2+7^x ... please explain >.<

OpenStudy (danjs):

y = (1/2) + 7^x

OpenStudy (danjs):

finding the inverse is the same as reflecting the function over the line y=x

OpenStudy (thatonegirl_):

yea

OpenStudy (danjs):

sqitch variables x and y, swap them around

OpenStudy (danjs):

x = (1/2) + 7^y

OpenStudy (thatonegirl_):

mhm.. then what??

OpenStudy (danjs):

solve for y x - (1/2) = 7^y

OpenStudy (thatonegirl_):

oh wait, its 1/(2+7^x)

OpenStudy (danjs):

\[\log _{7}(x-\frac{ 1 }{ 2 }) = y\]

OpenStudy (thatonegirl_):

\[y=1/(2+7^x)\]

OpenStudy (danjs):

oh

OpenStudy (thatonegirl_):

sorry

OpenStudy (danjs):

same process

OpenStudy (danjs):

x = 1 / (2 + 7^y)

OpenStudy (thatonegirl_):

wait why wouldn't it be 2x+14yx?

OpenStudy (danjs):

sorry here

OpenStudy (danjs):

x * (2 +7^y) = 1 2 +7^y = 1/x

OpenStudy (danjs):

7^y = 1/x - 2

OpenStudy (danjs):

\[y = \log _{7}(\frac{ 1 }{ x } -2)\]

OpenStudy (thatonegirl_):

THANK YOU SO MUCH

OpenStudy (danjs):

so y now equals the inverse function \[f ^{-1}(x)\]

OpenStudy (danjs):

no prob

OpenStudy (thatonegirl_):

:)